### v3

# DatAasee: A Metadata-Lake as Metadata Catalog for a Virtual Data-Lake

Abstract
Metadata management for distributed data sources is a long-standing but ever-growing problem. To counter this challenge in a research-data and library-oriented setting, this work constructs a data architecture, derived from the data-lake: the metadata-lake. A proof-of-concept implementation of this proposed metadata aggregator is presented, too, and also evaluated.
1 Introduction
In the shadows of data evolving to big data, metadata became big metadata. And due to metadata’s idiosyncrasy of describing other data, metadata management for big data warrants its own specialized discussion. In particular the inert dependency of metadata on its data needs to be accounted for, and leads to the driving conjecture behind this work and its object of study: existence, detail and accuracy of metadata determines discoverability, use and value of its associated data. To this end, a system for metadata management, which centralizes metadata from distributed data sources, is defined and its implementation described in the following.
This work focuses expressly on university libraries; and from a library perspective, the overall concept corresponds to a union catalog (Caplan 2003, Ch. 4) of publications and research data. Considering then that data operations (DataOps) is intertwined with metadata management (Reis and Housley 2022, Ch. 6; Strengholt 2023, Ch. 9), operationalizing research data founds upon enabling libraries to manage distributed research data’s metadata.
A starting point is the abstract data architecture: A data-lake is a central repository for the storage of (raw) data, typically together with at least some facility to ingest data (Mathis 2017; Serra 2024). The rationale for considering a data-lake is that the original raw (meta)data is still available for custom applications. A data-lake differs from a data warehouse in that it stores not only structured data but also semi-structured and unstructured data.
To make a data-lake navigable and prevent it from becoming a data swamp, a metadata catalog indexes the stored data by its metadata, as illustrated in the “Functional architecture of a data lake” (Laurent and Madera 2020, Ch. 1). Such metadata management in data lakes is furthermore characterized in Sawadogo et al. (2019) and references therein. Additionally, see Boukraa, Bala, and Rizzi (2024) for a review of metadata management in data-lakes.
This work discusses a special case of the data-lake: the metadata-lake [^2] One of the first mentions of the term “metadata-lake” is in Sankar (2021)., for which the admissible data is constrained to metadata encoded in plain-text. Given a setting where such a metadata-lake ingests metadata from various distributed data sources, it induces a virtual data-lake (see Section 4), due to the unified access of distributed data sources, while being its own metadata catalog, cf. Serra (2024), Ch. 6. A metadata-lake can also be seen as a data catalog, as it represents an inventory of available datasets. However, a data catalog typically builds a database from (global) metadata fields in linked databases (such as column descriptions), while the metadata-lake incorporates (descriptive) metadata from records in linked databases (such as bibliographic metadata). Another related metadata system is the data map (Rooney et al. 2021), which compared to the metadata-lake differs by focusing on tracking metadata changes, whereas the metadata-lake provides an aggregation of current metadata records.
The foundational data structure of the metadata-lake is the entity-relationship model, where entity attributes are handled as a “metadata repository supporting multiple formats without record conversion” (Zeng and Qin 2022, Ch. 8.5.2), while the entity relations are handled as “aggregation and enriched metadata records in a repository” (Zeng and Qin 2022, Ch. 8.5.3).
Beyond formally defining a metadata-lake, also the (open-source) implementation of the metadata-lake “DatAasee [^3] The Aasee is an artificial lake in Münster, Germany. (https://github.com/ulbmuenster/dataasee) is described. DatAasee is designed to centralize and unify bibliographic data and metadata of research data and aimed at universities and research institutes, with explicit application at the University and State Library of Münster.
Overall, this work is organized as follows: In Section 2 the potential use-cases for a metadata-lake are detailed. Section 3 reviews the types of metadata and derives a data model for metadata. Then in Section 4 an abstract definition for a metadata-lake is constructed, while in Section 5 the metadata-lake implementation DatAasee is described. Finally, Section 6 briefly evaluates DatAasee in terms of existing metadata system features and also proposes alternative criteria.
2 Use-Cases
Before describing the metadata-lake, six potential use-cases and applications for such a service are summarized. All of these were motive for investigating the metadata-lake concept and implementing the DatAasee metadata-lake.
2.1 Centralized Metadata
First and foremost, the purpose of the proposed metadata-lake is to centralize metadata from multiple sources, to provide a single point of access in a uniform format for university-wide (or company-wide) metadata and thus ultimately data, cf. Anderson (2022). This means distributed data is made available through its metadata by data-pipelining into a central repository, which is the central theme in data engineering (Reis and Housley 2022).
2.2 Cross-Discipline Discoverability
A related, but mainly academic motivation is interdisciplinary data discovery. Facilitating findability of datasets without knowledge of discipline-specific metadata formats requires cataloging the metadata in a common format. Aiding non-specialists, the metadata-lake partially transforms ingested metadata to a uniform schema which is discipline-agnostic.
2.3 Metadata Hub
While the previous applications focus on manual or semi-manual queries, a further use-case of the metadata-lake is to function as a data-hub for metadata-provisioning of other systems. In this setting all local444w.r.t. to the organizational unit the metadata-lake spans. metadata processing systems access metadata via the metadata-lake, which entails simplifications in their implementation and operation due to hiding heterogeneous source systems behind an intermediate metadata-layer.
2.4 FAIR-Compliant Repositories
The FAIR (Findable, Accessible, Interoperable, Reusable) principles (Wilkinson 2016) formalize good scientific practices discipline-agnostically. However, even if a particular research artifact is FAIR, research in general can only be FAIR if it, and particularly its metadata, is managed in a repository supporting FAIRness.
In Castro et al. (2022), a software reference architecture for general FAIR-compliant repositories is proposed. This architecture includes a so-called “metadata storage layer”, which could be realized by the metadata-lake proposed in this work [^5] Note that in Castro et al. (2022) the metadata storage layer contains a metadata-lake component, yet this work’s metadata-lake corresponds to the full “metadata storage layer”.. Accordingly, the metadata of research data objects from all “personal repositories” would be gathered in a metadata-lake, transformed and made accessible as well as associated research data retrievable through their identifiers.
2.5 Dataspace
A dataspace is a data management ansatz based on a set of data sources and relations among their contents, as well as a common interface for accessing the participating data source’s contents agnostic of their format (Curry 2020). By unifying data access to distributed data sources without integrating the data itself, conceptually a dataspace is a graph-based catalog for a virtual data-lake. In this sense, the herein proposed metadata-lake can instantiate such a dataspace.
2.6 Funding Compliance
Catalogs of metadata could become required by funding agencies, for example in Germany, as indicated in a government white paper about a planned research data law (Bundesministerium für Bildung und Forschung 2024), where the stated associated goal is improved research data discoverability. In this scenario public research institutes have to operate a metadata cataloging service, such as a metadata-lake, in order to comply with funding prerequisites.
3 Metadata
As this work is at the intersection of databases and research data or bibliographic metadata, the meaning of the term “metadata” needs to be defined in this context. Combining definitions from Pomerantz (2015) and Baca (2016), metadata is set to mean here:
Definition 1 (Metadata):
All statements about a (tangible or digital) information object are metadata.
For instance, metadata of a database is, in case of a relational database, a description of a table column. On the other hand, metadata about research data as well as bibliographic metadata is typically understood as a set of standardized key-value pairs describing the underlying data or published resource. The latter is the sole data of the metadata-lake and object of study in this work, which henceforth is referred to by the term “metadata” unless noted otherwise.
Following Sawadogo et al. (2019) and Ravat and Zhao (2019), metadata can be organized into three categories: intra-object metadata, inter-object metadata, and global metadata. Each of which is detailed next. Note, that intra-object metadata and inter-object metadata correspond to attributes and relations in the entity-relationship model (Zeng and Qin 2022, Ch. 4.1.1).
3.1 Intra-Object Metadata
Intra-object metadata refers to all literal values describing an underlying dataset. In the classic case of bibliographic metadata it is comprised of (among others) title, creator(s), publisher, publication date, type, identifier, etc. Given, that a (meta-)data processing service, such as a metadata-lake, also amounts metadata with respect to its data - the metadata, a more fine-grained classification of intra-object metadata types is advised. For the specific application of the metadata-lake the following classes are employed:
  • Descriptive Metadata, also known as business metadata, refers to information about the contents of a dataset, such as the bibliographic metadata mentioned above, and is denoted subsequently by $m_{\text{descriptive}}$.
  • Technical Metadata also known as structural metadata, describes properties of the metadata’s underlying dataset (i.e. a file or archive of files) such as location, format, size, or checksum, and is denoted by $m_{\text{technical}}$.
  • Processual Metadata, also known as operational metadata, encompasses information about the metadata record’s source, internal identifier, or creation and modification dates, and is symbolized by $m_{\text{processual}}$.
  • Administrative Metadata determines access eligibility for the metadata itself as well as the underlying dataset, and is denoted by $m_{\text{administrative}}$.
  • Social Metadata (Yassin 2022; Strengholt 2023, Ch. 9) refers to metadata related to discoverability and usage, such as keywords from controlled and uncontrolled vocabularies or number of views, and is denoted by $m_{\text{social}}$.
  • Raw Metadata refers to the originally ingested metadata from other data sources, justifies the data-lake aspect of the metadata-lake, and is denoted by $m_{\text{raw}}$.
Altogether, intra-object metadata $m_{\text{intra}}$ is the union of these classes:
$m_{\text{intra}}:=\{m_{\text{descriptive}},m_{\text{technical}},m_{\text{processual}},m_{\text{administrative}},m_{\text{social}},m_{\text{raw}}\}.$
Among these none is generally elevated, yet $m_{\text{raw}}$ is defining for this work. For instance, metadata management using intra-object metadata without $m_{\text{raw}}$ yields a data warehouse, while using intra-object metadata of exclusively $m_{\text{raw}}$ results in an unmanaged data-lake.
3.2 Inter-Object Metadata
While intra-object metadata collects literal data about a dataset, inter-object metadata describes the relation between underlying datasets. Following Sawadogo et al. (2019), relations between datasets are classified by:
  • Grouping relates datasets where either one is part of another or both are part of a third, and denoted by $m_{\text{grouping}}$.
  • Similarity relates datasets which are, the same work, manifestation, expression, or item using the terminology of FRBR (Functional Requirements for Bibliographic Records 2009); and is symbolized by $m_{\text{similarity}}$.
  • Parenthood relates datasets where one descends from the other, and is denoted by $m_{\text{parenthood}}$.
These interrelations between datasets are metadata which is represented by relations between metadata records, and jointly inter-object metadata is represented by $m_{\text{inter}}$ as union of these classes:
$m_{\text{inter}}:=\{m_{\text{grouping}},m_{\text{similarity}},m_{\text{parenthood}}\}.$
3.3 Global Metadata
Lastly, global metadata refers in this context to metadata about the metadata records, which corresponds to the metadata understanding of databases and potentially includes:
  • Database schema modeling the entities by defining and constraining attributes and relations.
  • Enumerated vocabularies listing standardized values used exclusively in certain attributes.
  • Context information, such as database comments, for instance in SQL by COMMENT ON DATABASE (PostgreSQL Global Development Group n. d.).
  • Labels and descriptions which provide human-readable phrases for attributes, relations and content details, with use, for example, in a frontend application.
Global metadata is rather implementation-specific and not of interest in the subsequent sections.
4 Metadata-Lake
This section provides a formal definition of the metadata-lake, which is based upon its foundational data architecture - the data-lake.
Definition 2 (Data-Lake)
A data-lake is a set of datasets.
Note, that the data may or may not be in some format (schema). Furthermore, since a set consists of different elements, it needs to be ensured that all elements are distinguishable (Maccioni and Torlone 2018), for example by a unique (but local, w.r.t. the metadata-lake) record identifier, which cannot be an identifier of the underlying dataset, i.e. a DOI (Digital Object Identifier), since a dataset maybe listed in multiple source systems.
Now, given the aim of this work, a metadata-lake cannot just be defined as a set of metadata records; beyond holding a record identifier and the original raw metadata, an internal metadata-lake format needs additional fields to provide a metadata catalog and fulfill the previously presented use-cases. Particularly, a field is needed to locate the metadata record’s underlying data if the metadata-lake is used as metadata catalog for a virtual data-lake.
Such a schema is derived using the metadata classification from Section 3. Overall, the metadata-lake is made up of a hierarchy of tuples:
Definition 3 (Intra-Object Attributes):
A set of ordered pairs $\{(K,V)_{i}\}$ is called intra-object attributes if:
  1. 1.
    $K_{i}\in L_{\text{intra}}$, with $L_{\text{intra}}$ being a finite set of labels,
  2. 2.
    there exists a map $f:\{V_{i}\}\to L_{\text{intra}}$ such that $\{K_{i}\}=\operatorname{image}(f)$.
Intra-object attributes conform to Definition 1, since one can implicitly insert the word “is” between a key and its value, and a statement is created of the form $K_{i}$ is $V_{i}$”. In practical technical terms intra-object metadata is a special case of a key-value document, typical for document databases.
If an internal identifier is used to ensure discriminability, it needs to be represented by an intra-object attribute, and similarly the raw metadata needs to be encoded by such a pair. Furthermore, an external identifier (i.e. a data location) should be included to actually span a virtual data-lake.
Since a common format for all metadata records is assumed, a space of all possible valid sets of such pairs can be called $M_{\text{intra}}$, and the associated set of labels $L_{\text{intra}}$ is abstractly defined in Section 3.1.
Definition 4 (Inter-Object Relations):
A set of ordered triples $\{(A,R,B)_{i}\}$ is called inter-object relations if:
  1. 1.
    $R_{i}\in L_{\text{inter}}$, with $L_{\text{inter}}$ being a finite set of labels.
  2. 2.
    $A_{i},B_{i}\in M_{\text{intra}}$
This is equivalent to the serialization of a directed labeled graph, where $M_{\text{intra}}$ is the set of nodes, the ordered pairs $(A_{i},B_{i})$ form the edges, and some map $g:\{(A_{i},B_{i})\}\to L_{\text{inter}}$ exists such that $\{R_{i}\}=\operatorname{image}(g)$.
If the labels conform to the (usual) constraint of being a predicate, then also by $A_{i}$ $R_{i}$ $B_{i}$” statements are given, cf. Semantic Triple (Wikipedia contributors 2024), and the inter-object relations fit the Metadata Definition 1.
Similarly, the space of all valid triples $M_{\text{inter}}$ is declared, and the set of labels $L_{\text{inter}}$ is abstractly defined in Section 3.2.
Definition 5 (Metadata Graph):
A graph $G$ where the vertices are a set of intra-object attributes $m_{\text{intra}}\subset M_{\text{intra}}$, and the edges are a set of inter-object relations $m_{\text{inter}}\subset M_{\text{inter}}$, connecting the former pairwise, is called (in this context) metadata graph,
$G:=(m_{\text{intra}},m_{\text{inter}}).$
With respect to Sawadogo et al. (2019), a metadata-lake can then be defined by:
Definition 6 (Metadata-Lake):
A metadata-lake is an ordered triple $(G,T,S)$, where $G$ is a metadata graph, $T$ is a set of transformations, and $S$ is a set of sources, such that $G:=\bigcup_{i=1}^{|S|}T_{f(i)}(S_{i})$ for some suitable map $f:\{1\dots|S|\}\to\{1\dots|T|\}$, with $|\cdot|$ denoting a set’s cardinality.
This definition is closely related to the data-lake model in Diamantini et al. (2018), Sec. 3, with the distinction that the sources $S$ are metadata sources instead of data sources ($D$), and the addition of transformations $T$. Also, the “metalake” from Strengholt (2023), Ch. 9 is a related concept.
This abstract definition of a metadata-lake induces already all components for a practical implementation. First, some kind of database [^6] The downstream access to the stored metadata can be realized as a semantic layer. to hold the metadata graph $G$, second, a processor providing transformations $T$ from source formats to the metadata-lake’s internal format, and third, a configuration holding an encoding of the sources $S$, so that the contents of $S$ can be transformed by $T$ and included into $G$. Figure 1 illustrates the full hierarchy of the metadata-lake.
Note, that the lake aspect of the metadata-lake is given by the raw (meta)data $m_{\text{raw}}$. This means, assuming a common minimal encoding of metadata in all sources and $M_{\text{intra}}$, and all transformations $T$ are the identity map, one would have a plain data-lake.
Figure 1: Abstract metadata-lake.
5 DatAasee
Figure 2: DatAasee metadata-lake.
Here the metadata-lake template derived in the previous section is instantiated. This means specifying an internal metadata format within a technical system. To this end, the following non-essential guidelines were predetermined:
  • A functional software architecture with minimal state.
  • Components are isolated in containers.
  • Components are implemented as declarative as possible.
  • Interaction is exclusively realized through an HTTP-API.
  • All API communication is in JSON encoding.
As stated earlier, the primary goal of the metadata-lake is centralizing metadata access, and since the metadata-lake is a data-lake and at the same time also a metadata catalog, the “Hub and Spoke Metadata Architecture” from Laurent and Madera (2020), Sec. 4.8.2 is abstractly employed. Its weakness of a customized standard meta model is mitigated by only partially parsing ingested metadata records and thus reducing the extent of overall load.
In a top-down approach the metadata-lake architecture is subsequently staked out in ever more detail. For an overview of the practical data model (see also Section 5.2.1.1) and the application of the metadata-lake, see Figure 2 and Figure 3.
Figure 3: DatAasee outward architecture.
5.1 Data Architecture
The most foundational decision about building a (metadata management) service is its data architecture. Fundamentally, the data architecture is fixed already as data-lake, but only to the degree of Definition 2. Implementing a data-lake requires additional practical architectural considerations (Hai et al. 2023).
5.1.1 Data-Lake Architecture
A process pattern that is closely related to data-lakes is ELT (Extract-Load-Transform), which summarizes a process that extracts data from a source system, loads it as-is into the destination system, and transforms it on-demand. On the other hand, the corresponding process pattern in data warehouses is ETL (Extract-Transform-Load), where data is extracted from a source system, transformed according to the destination system’s schema, and loaded into it. The metadata-lake stores raw metadata like a data-lake, as well as transformed data like a data warehouse due to its metadata catalog aspect. Hence, the EtLT (Extract-transform-Load-Transform) process patterns from Densmore (2021) is effectively used. This means the raw metadata is partially transformed to match the metadata-lake’s catalog schema, while retaining the original raw metadata which is then transformed to its (external) application upon request.
Given typical data-lake architectures (Hlupić et al. 2022), the metadata-lake corresponds to the two-layered architecture with a transient landing zone and a second zone that divergingly not only holds the raw (meta)data, but jointly raw and refined metadata. This directly results from the EtLT process choice.
5.2 Software Architecture
Figure 4: DatAasee inward architecture.
The foregoing data architecture is implemented following a classic three-tier software architecture (Manuel and AlGhamdi 2003), with the typical tiers:
  • Data-tier: A database management system (DBMS),
  • Application-tier: A stateless processing system encoding the domain logic,
  • Presentation-tier: A semantic layer as interface for users or clients,
which are specified next; see also Figure 4.
5.2.1 Data-Tier
Similar to Prabhune et al. (2016), DatAasee is based on a NoSQL database system, which stores all types of metadata summarized in Section 3. Here the inter-object relations suggest support of a graph model (Ravat and Zhao 2019), while the intra-object attributes imply a document (or relational) model. A property-graph database matches both needs by using documents as nodes and edges. Thus, the proposed metadata model translates as follows: The intra-object attributes correspond to the nodes (vertices), and the inter-object relations are realized by edges, cf. entity-relationship model.
Practically, the DBMS ArcadeDB (Garulli et al. n. d.) is employed in DatAasee. ArcadeDB is a multi-model NoSQL database featuring a hierarchy of models: An overarching graph model with a document model for nodes and edges, and a key-value model for document contents, resulting in a property-graph.
A specialty of ArcadeDB is its polyglotism: While the native query language is a dialect of SQL, Gremlin and Cypher are also supported. So, the commonplace SQL eases entry, while specific graph query-languages are available, too.
5.2.1.1 Data Model
The data model of the DatAasee data-tier follows the metadata-lake schema proposed in Section 3 and Section 4: Intra-object metadata is stored in a vertex, and inter-object metadata is encoded by edges. The vertices consist of descriptive, technical, processual, social, and raw metadata, where the descriptive metadata key-value pairs are derived from the DataCite 4.5 metadata kernel (DataCite Metadata Working Group 2024). Furthermore, administrative metadata is omitted, since all metadata records are unrestrictedly available to all users for reading (license and rights fields are kept as descriptive metadata).
In terms of a relational database, the metadata vertex schema (intra-object attributes) would correspond to a wide table with a star schema for enumerable dimensions. Edge types are a subset of the DataCite item relations together with a subset of the OpenWEMI (Dublin Core Metadata Initiative 2024) entity relations, of which the latter encode the FRBR. For details about the database schema, see the Appendix.
In the metadata-lake, the two mandatory intra-object attributes are the record identifier and raw metadata. While the former is computed as a xxh64 hash from a subset of descriptive and processual metadata and then base64url encoded, the latter is stored as a plain string, since only textual metadata was assumed.
5.2.1.2 Alternative Implementations
The property-graph-based implementation chosen for DatAasee is not imperative. Likewise, a relational or knowledge-graph-based data-tier could have been used. In case of a relational database, the inter-object metadata would be emulated by a single or multiple junction tables, while in a knowledge graph, the intra-object metadata can be constructed via statements with literal objects. Thus, if implementation constraints or additional requirements demand it, alternative database types are compatible. However, in the author’s opinion a property graph is the ideal choice for the proposed metadata-lake.
5.2.2 Application-Tier
Basically, the domain logic consists of two components, the import of metadata on the one hand, and the provisioning of it on the other. The import from source systems includes the request-driven transport, parsing, (partial) transformation, data-cleaning, metadata quality evaluation and storing in the data-tier. Provisioning happens by listening for presentation-tier requests, translating those to data-tier queries, and re-translating the results to a domain response, akin to a semantic layer. Requests for a single record by its identifier or search by filter, full-text or custom query in a compatible query-language are supported.
5.2.2.1 Data Pipelines
The source and transformation components of the metadata-lake, with itself as implicit destination, form data pipelines. A data pipeline needs to be specialized by a configuration to be able to admit updates or additions during operation. Technically, four (metadata) items are minimally needed to specify a data source and transformation for DatAasee:
  • Location: is a string that translates to a resolvable address, i.e.: URL.
  • Protocol: specifies sequence and content of messages, i.e.: OAI-PMH.
  • Encoding: fixes the manner of metadata payload formulation, i.e.: XML.
  • Format: then describes how the content is arranged, i.e.: DataCite.
Location and protocol are required to extract the metadata from its source, while encoding and format are essential to (partially) transform the metadata into the native metadata-lake schema. Potentially, also credentials (i.e.: username/password, access-/secret-key) maybe necessary to gain access to a source. Non-technically, DatAasee requires a data-steward to be named for every ingested source, as well as the rights (license) under which the metadata of the source are ingested. Beyond these, an options field is available to pass protocol specific configuration, i.e. selective harvesting for OAI-PMH.
Currently, DatAasee supports the OAI-PMH, S3, and GET protocols, as well as DataCite, DublinCore, LIDO, MARC, and MODS formats [^7] See openarchives.org/pmh, docs.aws.amazon.com/s3, httpwg.org/specs/rfc9110, dublincore.org, lido-schema.org, loc.gov/standards/marcxml, and loc.gov/standards/mods. Encoding-wise XML is implicitly assumed, but is principally extensible to JSON, CSV or others. Additionally, the contents of another DatAasee instance can be ingested.
For exporting metadata, DatAasee provides its native format which is derived from DataCite, as well as actual DataCite, and BibJSON (MacGillivray and Pitman n. d.) - a JSON representation of BibTEX.
5.2.3 Presentation-Tier
As an API-based web-service, cf. Data Commons (Grossman et al. 2016), the primary front-end of the DatAasee metadata-lake is a REST-like (REpresentational State Transfer) API (Application Programming Interface) with CQRS (Command-Query-Responsibility-Segregation) aspects over HTTP. While the ingest encoding is XML, the HTTP-API only accepts and responds exclusively in JSON. As a secondary effect, DatAasee can also serve as translator from legacy XML sources to JSON. The response content is formatted following the JSON-API (Katz et al. 2022) standard. And to establish self-documenting messages, each response contains a link to its content’s associated JSON Schema, such that the wrapping JSON-API and its payload data are understandable. Furthermore, this API is documented via an OpenAPI (Miller et al. 2020) document which is provided by the API itself. Since all messages to and from the DatAasee API server are in JSON encoding, for all requests and responses JSON Schemas are provided through the API itself.
5.2.3.1 Web-Frontend
In addition, to the primary HTTP-API interface, a secondary prototype graphical user interface for the metadata-lake in the form a web application is available, which in turn exclusively uses the HTTP-API. While optional and fully decoupled from the metadata-lake proper, and beyond being a demonstrator facilitating exploration and review, it serves the purposes of:
  • template implementation of an exemplary graphical user interface;
  • living documentation, for a hands-on interaction with the metadata-lake;
  • more comfortable manual testing, and abstracting lower-level formatting.
6 Preliminary Evaluation
There are various evaluation criteria for metadata systems of data-lakes, such as the functional features: Semantic Enrichment, Data Indexing, Link Generation, Data Polymorphism, Data Versioning, and Usage Tracking from Sawadogo et al. (2019), the generic features: Metadata Flexibility, Granular Levels, and Multiple Zones from Eichler et al. (2020), or the semantic features: Semantic Labels, Semantic Relationships, Central vs External Graph, Metadata Interoperability, Initial Automatic Labeling, Technical Abstraction, and Open-Source from Hoseini, Theissen-Lipp, and Quix (2023). Notably, these features are primarily concerned with the internal organization of the metadata management. In Table 1, a brief evaluation of DatAasee in terms of the aforementioned features is summarized; for a description of these features see the respective references.
Table 1: Overview of DatAasee Features
Feature DatAasee Comment
Semantic Enrichment Yes Schema includes categorization
Data Indexing Yes Data-tier runs (full-text) indices
Link Generation Yes Based on relations and identifiers
Data Polymorphism Yes Raw and refined metadata stored
Data Versioning No Only if tracked via relations
Usage Tracking Yes Via social metadata
Metadata Flexibility Yes Uses document vertices and edges
Granular Levels No Only via grouping relations
Multiple Zones Yes Raw and refined zone
Semantic Labels No Only in terms of DataCite relations
Semantic Relationships Yes Via DataCite relations
Central vs External Graph No Federation is not yet implemented
Metadata Interoperability Yes Exportable as JSON
Initial Automatic Labeling Yes Final step of ingest process
Technical Abstraction No Advanced use needs query languages
Compatibility No Only with DataCite and OpenWEMI
Open-Source Yes MIT-licensed
6.1 FAIR as Alternative Features
An alternative set of features for the evaluation of metadata layers or metadata systems can be derived from the FAIR principles [^8] See also: https://www.go-fair.org/fair-principles. Those are not meant to be added, but rather as an alternative, since their coverage of underlying qualities overlaps with the features above. In comparison, the following criteria are concerned with the outward functionality instead of internal structure:
  • Findability is given by the exposed search modes, such as full-text (F1), filters or facets (F2), or via query languages (F3).
  • Accessibility is provided by an API, which should follow a standard (A1), be self-explaining (A2), and documented (A3).
  • Interoperability refers to flexibility in protocols (I1), encodings (I2) and formats (I3) from which can be imported and to which can be exported.
  • Reusability is enabled by open source code licensing (R1), multi-layered documentation (R2) and hardware/software compatibility (R3).
All of the above are checked by DatAasee.
7 Summary
A foundational attribute for research data, research software, or publications to be useful is first and foremost their availability (Fehr et al. 2016). So, a data-lake - or metadata-lake - with its priority of load before transform (in the context of ELT) qualifies for the task of research metadata management. The metadata cataloging aspect of the metadata-lake then makes it a suitable metadata layer for a virtual data-lake, dataspace, or FAIR-compliant repository.
DatAasee is an open-source implementation of the proposed metadata-lake architecture, and aims to become the metadata-layer in academic institutions such as universities and research institutes. Particularly, to automate metadata handling and improve research data discoverability in university libraries, academic libraries, scientific libraries or research libraries.
Code Availability Statement
The source code for DatAasee (version 0.5) is available under MIT-license at:
Acknowledgement
The author thanks Holger Przibytzin, Marc Schutzeichel, and Jan-Erik Stange for providing valuable feedback on this manuscript.
Appendix
Below is the DatAasee data-tier database schema, encoded in yasql [^9] See: github.com/aryelgois/yasql, visualized: